A Gardener has 60 feet of fencing to enclose a rectangular garden and the two foot wide border surrounding it.
If The length of the garden is twice its width, what are its dimensions?

Answer by checkley71(8403)

GARDEN SIDES=X&2X
GARDEN PLUS BORDER=(X+4)&(2X+4)
PERIMETER=2(X+4)+2(2X+4)=60
P=2X+8+4X+8=60
P=6X+16=60
6X=60-16
6X=44
X=44/6
X=7.33 LENGTH OF THE GARDEN
2*7.33=14.66 WIDTH OF THE GARDEN.
PROOF
2(7.33+4)+2(14.67+4)=60
2*11.33+2*18.67=60
22.66+37.24=60
60=60